What are the coordinates of the point 2/5 of the way from h to e on the grid
1
Expert's answer
2020-10-21T17:24:09-0400

let H coordinates be ("x"1,y1)

let E coordinates be ("x"2,y2)


 lets find the coordinates of the point 2/5 of the way from h to e on the grid

A ("x" , y)

using section formula

"x" = "x"1 + "\frac{a}{a+b}" ("x"2 - "x"1) and y = y1 + "\frac{a}{a+b}" (y2 - y1)

a=2 and b=5

"x" = "x"1 + "\frac{2}{2+5}" ("x"2 - "x"1) and y = y1 + "\frac{2}{2+5}" (y2 - y1)


"x" = "x"1 + "\frac{2(x_2-x_1)}{7}" and y =y1 + "\frac{2(y_2-y_1)}{7}"


"x" = "x"1 + "\frac{2x_2-2x_1}{7}" and y = y1 + "\frac{2y_2-2y_1}{7}"


"x" = "\frac{7x_1+2x_2-2x_!}{7}" and y = "\frac{7y_1+2y_2-2y_1}{7}"


"x" = "\frac{5x_1+2y_2}{7}" and y = "\frac{5y_1+2y_2}{7}"


these are the new coordinates


A( "\frac{5x_1+2x_2}{7}" , "\frac{5y_1+2y_2}{7}" )